You are considering the purchase of a stock that is not currently listed on an active secondary market. On the recent financial statement, the book value of equity is reported as $36,425,000 and there are 6,257,000 shares outstanding. You plan on calculating the Market‐to‐Book Ratio for a comparable company whose shares are publically traded, and are currently selling for $27.93. This firm’s book value of equity is $139,843,000 and it has 27,826,000 shares outstanding. Based on this comparable Market‐to‐Book Ratio, how much should you be willing to pay for the shares of your firm of interest?
Answer #1

\textup{Market to Book Ratio} =\frac{ \textup{Market Capitalization} }{\textup{Total Book Value}}

Let us first calculate the Market to Book Ratio so that we can use that ratio to find the acceptable price of our non listed stock.

Therefore, Market to Book Ratio of the comparable company =

=\frac{ \textup{27.93 *27,826,000 } }{\textup{139,843,000}}

= 5.56

Using this information we can find the acceptable price of our target company.

Let us use the same formula:

\textup{Market to Book Ratio } = \frac{\textup{Market Capitalization}}{\textup{Total Book Value}}

Let the acceptable price of share be "x"

5.56 = \frac{\textup{6,257,000}*x}{\textup{36,425,000}}

\therefore x = 32.35

Therefore, the amount willing to pay for an unlisted share of the company is $32.35.